# A Genealogy of Multi-Sensor Foundation Models in Remote Sensing

### v1

@@ Abstract.
Foundation models have garnered increasing attention for representation learning in remote sensing, primarily adopting approaches that have demonstrated success in computer vision with minimal domain-specific modification. However, the development and application of foundation models in this field are still burgeoning, as there are a variety of competing approaches that each come with significant benefits and drawbacks. This paper examines these approaches along with their roots in the computer vision field in order to characterize potential advantages and pitfalls while outlining future directions to further improve remote sensing-specific foundation models. We discuss the quality of the learned representations and methods to alleviate the need for massive compute resources. We place emphasis on the multi-sensor aspect of Earth observations, and the extent to which existing approaches leverage multiple sensors in training foundation models in relation to multi-modal foundation models. Finally, we identify opportunities for further harnessing the vast amounts of unlabeled, seasonal, and multi-sensor remote sensing observations.
@@ 1. Introduction
With the recent successes of Natural Language Processing (NLP) frameworks like Generative Pre-trained Transformers (GPT) (OpenAI et al., 2024) and Bi-directional Encoder Representations from Transformers (BERT) (Koroteev, 2021), the development and application of machine learning is shifting away from focusing on transfer learning, which utilized traditional supervised learning on a large, labeled dataset to learn representations of that data to apply to other related tasks. Instead, Self-Supervised Learning (SSL), which attempts to learn general-purpose representations of data, known as encodings, has risen in popularity, showing considerable improvement in the quality of learned representations (Chen et al., 2024). These models have a number of advantages over their predecessors, particularly in their ability to leverage vast amounts of unlabelled data. As datasets have grown larger, with some datasets now containing billions of datapoints (Dehghani et al., 2023), the process of labeling data for supervised learning has proven to be a costly bottleneck for improving model performance. Learning encodings through SSL has allowed researchers to fine-tune models for specific downstream tasks on significantly smaller labeled datasets in order to achieve state of the art (SOTA) performance on a single task. These generalizable representations can be utilized for any number of downstream tasks within a domain. Decoupling the process of learning encodings from the process of learning downstream tasks enables downstream users to leverage high performance machine learning models and minimizes the amount of labeled training samples needed to learn specific tasks.
In the wake of NLP success, the computer vision (CV) field swiftly delved into SSL with models such as BYOL (Grill et al., 2020), SimCLR (Chen et al., 2020), MoCo (He et al., 2020), and DINO (Caron et al., 2021), all of which use unique approaches to create self-supervised tasks and objectives for natural imagery data. These methodologies have recently started to percolate into the field of remote sensing (RS), where the sheer amount of unlabelled data from satellite imagery makes the field ripe for the development of foundation models. The importance of enabling training with fewer labeled samples for environmental monitoring in the face of climate change, such as monitoring or forecasting wildfires and flooding, means that there is imminent use for these generalized representations for downstream tasks. Thus far, the self-supervised methods for foundation models in RS build heavily upon CV methods and fall broadly into one of three categories:
@@
Masked auto-encoding
Most of these approaches have demonstrated significant promise in the RS domain but have largely remained limited to applying the same CV frameworks to satellite imagery and other remote sensing datasets with few to no methodology changes (Rolf et al., 2024). On the other hand, recent research demonstrates how satellite data has unique properties that are distinct from natural imagery, such as the logarithmic distribution of objects and the range of the electromagnetic spectrum captured within satellite images (Rolf et al., 2024). These properties indicate that satellite imagery is a distinct modality and would therefore benefit from modality-specific frameworks and self-supervised learning tasks. Moreover, the variety of different satellite and ground-level sensors available allows for the development of highly specialized multi-modal foundation models.
Recent papers in the field have surveyed available foundation models in RS. Zhou explored the high potential of applied Vision Language Models (VLMS) while highlighting their few shot capabilities in downstream tasks (Zhou et al., 2024). Lu examined the model backbones used in foundation models, taking care to touch on ResNet (He et al., 2015) and Transformer (Vaswani et al., 2023) backbones, and compiled results for RS foundation models across a wide range of diverse downstream tasks (Lu et al., 2025). Chen inspected how models can adapt to utilize non-raster based data (Chen et al., 2024), while Zhang examined the singular task of building foundation models specifically for urban environments (Zhang et al., 2024), and Mai created a list of traits desirable in RS foundation models (Mai et al., 2023a). While these works provide a timely review of important components in RS foundation models, none of them focus on understanding the origins of RS SSL techniques and the modifications that have been developed for handling RS data, or those modifications that intend to alleviate the need for substantial computational resources. These modifications are critical to model performance and understanding them is necessary for the development of future multi-modal RS foundation models. In this paper, we contribute to the growing body of research on foundation models in remote sensing by doing the following:
@@
We survey recent foundation models developed in the field, categorized by their learning objective, approach and compute requirements, and discuss their specific adaptations for RS observations.
We identify gaps, and lay out research opportunities for foundation models that specifically consider the multi-sensor and multi-modal nature of RS observations.
A variety of satellite images capturing Boston, MA in the month of October, 2020. Images with VV and VH polarizations captured by Synthetic Aperture Radar (SAR) are in black and white and do a good job of capturing water and urban structures. A false color image of SAR combines the different polarizations to provide a sharper, 3-channel image. There is a standard RGB optical image, and a false color image with SWIR bands and the blue light band that is used to highlight geological features. Then, there is a depiction of NDVI, where dark green highlights healthy vegetation, dark blue highlights no vegetation, and white indicates sparse or unhealthy vegetation.
@@ 2. Data for Training Remote Sensing Foundation Models
The CV field has primarily been concerned with natural images, which only sense the red, green, and blue wavelengths of the visible spectrum, and are typically captured at ground level. Some models, such as CLIP (Radford et al., 2021), FLIP (Jia et al., 2021), VisualBERT (Li et al., 2019), ViLBERT (Lu et al., 2019), FLAVA (Singh et al., 2022), and DeepSeek-VL2 (Wu et al., 2024) are VLMs that also utilize text as an additional mode of data. However, most of the RS foundation model frameworks borrowed from CV concern themselves with only the image modality. Within the RS domain, the variety of unique sensors available ranges across a number instruments aboard different satellites. The data modes detailed below are available at mass scale and readily accessible for training foundation models, which is our focus in this manuscript, rather than providing a full catalogue of remote sensing or geospatial data. We will keep discussions focused on sensors with high resolution (e.g., less than 40m), which are also the focus of current efforts in benchmarking datasets (Lacoste et al., 2023) and training foundation models (Zhou et al., 2024).
@@ 2.1. Optical Imagery
Optical images capture electromagnetic wavelengths in the visible and near visible bands of light. These are recorded via passive sensors, which simply measure the electromagnetic radiation originating from the sun that is reflected off of objects. This means that passive sensors are more energy efficient than active sensors, but they are affected by light and weather conditions. These sensors can only capture imagery during the daytime, and can be obscured by atmospheric phenomena like cloud cover or haze, making their use limited in hurricanes, storms or active floods under overcast conditions.
Optical images are often acquired from different satellites, with several openly available freely for mass download, such as MODIS (Justice et al., 2002), the Landsat (Wulder et al., 2022) family, and Sentinel-2 (Spoto et al., 2012), with the latter two offering high resolution. Each satellite contains a sensor with different spatial and spectral resolutions, but each is broadly designed to capture the visible, near-infrared, and short-wave infrared portions of the electromagnetic spectrum. In the case of the Sentinel-2 satellite, there are 13 distinct bands that are captured at resolutions ranging from 10m/pixel to 60m/pixel, and can be broken down into five groups based on usage as follows (Wang et al., 2016):
@@
B5-8 refer to VNIR (Visible and Near Infrared) light, and are typically used to measure vegetation health. B8a is a Narrow NIR band for moisture-based vegetation health monitoring.
B9, still in NIR spectrum, focuses on water vapor absorption and is used to measure atmospheric moisture presence, and paired with B8A to derive total precipitable water.
B10-B11 are SWIR (Short Wave Infrared) bands, with B10 for cirrus cloud detection, B11 for soil and vegetation moisture detection and snow/ice vs. cloud differentiation, and B12 is often used with B2 to measure geological features and mapping burned areas.
This range of spectrum and information captured by Sentinel-2 is an example of the wealth of information is captured by remote sensing optical products that go beyond RGB. However, CV models are designed to operate with three input channels of RGB of ‘natural’ images (obtained by regular cameras and widely available on the internet). In order to leverage the highly sophisticated pre-trained weights that are available for CV models such as the ubiquitous ResNet50 or ViT16, some RS models, such as SeCo (Mañas et al., 2021) or MMCPP (Chen et al., 2023), only utilize the visible bands of light from MS satellite images. RS models also rely on RGB images taken at ground level (Cepeda et al., 2023), or from some sort of aerial platform, such as a drone or helicopter (Liu et al., 2024). While RS models trained purely on RGB can achieve SOTA performance in certain settings (Mañas et al., 2021), deep learning models in RS that only train on RGB data have an information disadvantage compared to models that train on all bands and can struggle to perform on a number of downstream tasks in RS as a result (Lacoste et al., 2023). The spectral range of interest and the specific choice of bands on a satellite are the result of decades of experience, in-lab experiments and consensus-building efforts to ensure satellites are equipped with bands that can assist in target-specific tasks. Figure 1(f) highlights the information lost when solely examining RGB imagery. For instance, infrared bands are crucial for monitoring vegetation (Neinavaz et al., 2021), as seen in figure 1(f), while NIR bands in confluence with blue and green bands can help with water quality analysis (Cillero Castro et al., 2020). Shortwave infrared (SWIR) can be used for measuring soil and vegetation moisture, and is depicted in figure 1(e).
While increasingly, foundation models leverage most bands present in optical imagery such as Sentinel-2, they typically do so by a simple manipulation of input channel dimensions to the image encoder of the model (Wang et al., 2022). The optimal architecture required for remote sensing architectures is not well-defined yet (Rolf et al., 2024).
@@ 2.2. Synthetic Aperture Radar
Synthetic Aperture Radar (SAR) is an active remote sensor that emits energy in microwave frequencies towards the earth and then captures the amount of energy reflected back. Capturing the image from a moving sensor (in this case, a satellite) allows for capturing the image through a large ’synthetic’ aperture, compared to the much smaller physical sensor aperture, resulting in a much higher resolution than other sensors that monitor similar frequencies (Brown and Porcello, 1969). Each SAR sensor operates in a specific wavelength, referred to as bands, which serve different uses (Tsokas et al., 2022). Larger wavelengths provide more penetration through obstruction, such as vegetation, but result in lower resolution. One of the most commonly used bands is the C-band (7.5-3.8 cm wavelength) (Tsokas et al., 2022), which is currently captured by Sentinel-1 (Torres et al., 2012), RADARSAT-1 (Mahmood, 2014), RADARSAT-2 (Morena et al., 2004), RADARSAT Constellation Mission (RCM) (Thompson*, 2015) satellites, and made available publicly, and is useful for global mapping and change detection, as its larger wavelength provides moderate penetration still at relatively high spatial resolution (e.g., 40m). However, there are larger wavelengths available, such as L-band, which is captured by AirSAR (Hoekman and Bouman, 1993), PALSAR (Rosenqvist et al., 2007), and the soon-to-be-launched NISAR (Chapman and Rosen, 2024). L-band is often used for biomass and vegetation mapping (Haldar et al., 2012). There is also X-band, whose wavelength is smaller than C-band and often used for high resolution monitoring of fine surface features, such as vegetation structure, and urban infrastructure. (Solari et al., 2017). X-band is obtained by platforms such as Capella (Castelletti et al., 2021), ICEYE (Ignatenko et al., 2020), TerraSAR-X (Werninghaus and Buckreuss, 2010), and TanDEM-X (Krieger et al., 2007). Different wavelengths of SAR capture different information, and different wavelengths often complement each other, if available for the same region and time.
@@
SAR has another distinct characteristic from optical imagery: SAR takes advantage of materials’ innate dielectric properties; every material has a dielectric constant that describes how strongly it interacts with electromagnetic waves. Materials with a high dielectric constant, such as water, reflects microwave energy back more strongly, allowing it to be picked up more easily by SAR (Barowski et al., 2019). This can prove useful for applications such as mapping waterways in a region that might be otherwise undetected by optical sensors. As seen in 1(f), SAR imagery clearly highlights water features that optical sensors struggle to pick up. SAR is also an all-sky sensor, meaning that it operates regardless of light conditions, and its wavelength range is large enough that it penetrates cloud cover, which helps provide continuous spatial coverage in geographic areas that might otherwise be occluded to optical satellite sensors due to cloud cover (Brown and Porcello, 1969), or polar regions that lack sunlight for half of a year.
@@
SAR is predated by passive microwave sensing, which monitors similar microwave frequencies in order to pierce cloud cover but does not emit a signal of its own in order to do so. As a result, passive microwave sensing offers much lower resolution compared to SAR. Sensors such as AMSR (Tachi et al., 1989), AMSR-2 (Alsweiss et al., 2015), SMAP (Entekhabi et al., 2010), and GPM (Munchak et al., 2020) all passively monitor microwave signals emitted from the Earth. By comparison, SAR actively beams energy down to Earth’s surface and records the response, which allows for multiple polarizations at transmission and reception. Polarization describes the orientation of the plane that the electromagnetic wave oscillates on. The shorthand for polarization is the orientation of the emission, which is either vertical (V) or horizontal (H), followed by the orientation of the reception, which will also be either vertical or horizontal. Each polarization results in the waves scattering differently upon reflection, enabling certain polarizations to be particularly effective at picking up certain types of surface features. For example, the VV polarization is good at picking up rough surface scattering, typically caused by roughened water or vertical surfaces. HV or VH polarizations are good at picking up volume scattering, where the signal bounces around a number of times in a given volume before being broadcast back out to space; this is typically caused by forest canopies and vegetation. We see in figure 1(a) and figure 1(b) the the VH polarization provides more fine-grained details in a variety of forested areas than the VV polarization, where those areas are whited out. HH polarization is good at picking up horizontal surfaces, such as smooth ice or calm water.
@@
Compared to optical imagery, SAR captures distinct and complementary information, despite it also being stored as a raster (gridded) data type. The different frequencies of SAR, the different polarizations, and the volume scattering aspect make it stand apart from optical. For instance, surface reflectance of sea ice will appear as a bright surface in optical imagery, regardless of its age (which is a proxy for its thickness and level of hazard it poses to marine navigation). However, as sea ice ages (from thinner first-year ice to thicker multi-year ice), its salinity changes, and because of SAR’s volume scattering (caused by salt content and air bubbles in ice), the appearance of thick ice changes in SAR imagery compared to younger, thinner ice (Pires de Lima et al., 2023). This is the primary reason sea ice analysts primarily rely on SAR imagery in making sea ice charts, and use optical only as a secondary reference. As we demonstrate later in this manuscript, this complementary aspect of SAR imagery (despite it being stored as an image) motivates incorporating this multi-sensor information into RS foundation models.

### v2

# A Genealogy of Foundation Models in Remote Sensing

@@ Abstract.
Foundation models have garnered increasing attention for representation learning in remote sensing. Many such foundation models adopt approaches that have demonstrated success in computer vision with minimal domain-specific modification. However, the development and application of foundation models in this field are still burgeoning, as there are a variety of competing approaches for how to most effectively leverage remotely sensed data. This paper examines these approaches, along with their roots in the computer vision field. This is done to characterize potential advantages and pitfalls, while outlining future directions to further improve remote sensing-specific foundation models. We discuss the quality of the learned representations and methods to alleviate the need for massive compute resources. We first examine single-sensor remote foundation models to introduce concepts and provide context, and then place emphasis on incorporating the multi-sensor aspect of Earth observations into foundation models. In particular, we explore the extent to which existing approaches leverage multiple sensors in training foundation models in relation to multi-modal foundation models. Finally, we identify opportunities for further harnessing the vast amounts of unlabeled, seasonal, and multi-sensor remote sensing observations.
@@ 1. Introduction
With the recent successes of Natural Language Processing (NLP) frameworks like Generative Pre-trained Transformers (GPT) (OpenAI et al., 2024) and Bi-directional Encoder Representations from Transformers (BERT) (Devlin et al., 2019), the development and application of machine learning is shifting away from focusing on transfer learning. Transfer learning uses traditional supervised learning on a large, labeled dataset to learn representations of that data to apply to other related tasks. Instead, Self-Supervised Learning (SSL), which learns general-purpose representations of data without the need for external labels, has risen in popularity (Chen et al., 2024). As datasets have grown larger, with some datasets now containing billions of data points (Dehghani et al., 2023), the process of labeling data for supervised learning has proven to be a costly bottleneck for improving model performance. SSL circumvents this issue with the use of pretext tasks: auxiliary learning objectives designed to leverage the natural structure of unlabeled data in order to learn patterns or relationships within the data. Researchers can then fine-tune these pre-trained models for specific downstream tasks on significantly smaller labeled datasets. Fine-tuning enables downstream users to achieve state-of-the-art (SOTA) performance on a specific task while incurring significantly lower costs than training an equivalent model from scratch (Anisuzzaman et al., 2025). Thus, effective SSL training democratizes access to high performance machine learning models.
In the wake of NLP success, the computer vision (CV) field swiftly delved into SSL with foundation models such as BYOL (Grill et al., 2020), SimCLR (Chen et al., 2020), MoCo (He et al., 2020), and DINO (Caron et al., 2021), all of which use unique approaches to create self-supervised tasks and objectives for natural imagery data. These methodologies have recently started to percolate into the field of remote sensing (RS), where the sheer amount of unlabeled data from satellite imagery makes the field ripe for the development of foundation models. Satellite imagery can be used for environmental monitoring tasks, such as monitoring or forecasting wildfires. Climate change’s impact on such events means that there is imminent use for these generalized representations for downstream tasks.
Thus far, the self-supervised methods for foundation models in RS build heavily upon CV methods and fall broadly into one of four categories:
@@
Contrastive learning via redundancy reduction
Masked image modeling
Most of these approaches have demonstrated significant promise in the RS domain but have largely remained limited to applying the same CV frameworks to satellite imagery and other remote sensing datasets with few to no methodology changes (Rolf et al., 2024). However, recent research demonstrates how satellite data has unique properties that are distinct from natural imagery, such as the logarithmic distribution of objects and the range of the electromagnetic spectrum captured within satellite images (Rolf et al., 2024). These properties indicate that satellite imagery is a distinct modality and would therefore benefit from modality-specific frameworks and self-supervised learning tasks. Moreover, the variety of different satellite and ground-level sensors available allows for the development of highly specialized multi-modal foundation models. In this paper, we contribute to the growing body of research on foundation models in remote sensing by doing the following:
@@
We survey recent foundation models developed in the field, categorized by their learning objective, and discuss their specific adaptations for RS observations.
We identify gaps and outline research opportunities for RS foundation models. We specifically consider the multi-sensor and multi-modal nature of RS observations, and how to effectively leverage such data at scale.
2. Prior Work
Recent papers in the field have surveyed available foundation models in RS. Zhou explored the high potential of applied Vision Language Models (VLMS) while highlighting their few shot capabilities in downstream tasks (Zhou et al., 2024). Chen inspected how models can adapt to utilize non-raster based data (Chen et al., 2024), such as poly-lines, to represent transportation networks. Zhang examined the singular task of building foundation models specifically for urban environments (Zhang et al., 2024). Mai created a ’wish list’ of desirable traits in RS foundation models, with rationale for each wish, and discussed where current foundation models exceed or fall short of these goals (Mai et al., 2023a). Lu categorized the model backbones, training datasets, image resolution, and over-arching pre-training methods for existing RS foundation models (Lu et al., 2025). Lu also documented the number of parameters for each model and compiled results across a wide range of diverse downstream tasks (Lu et al., 2025).
While these works provide a timely review of important components in RS foundation models, none of them focus on the specific adaptations that have been made to SSL approaches for working with remote sensing data. Works such as (Zhu et al., 2024) might note that ideal RS foundation models should factor in temporal features, but do not spend time exploring the ins-and-outs of how current models do so. Surveys such as (Lu et al., 2025) comprehensively document the different datasets, backbones, and broad SSL tasks used by RS foundation models. However, they do not delineate the differences in models past this point, preferring to focus instead on model performance. While comprehensive benchmarking is helpful, differences in models’ learning objectives such as the use of temporally positive pairs, adaptation to multi-scale view, or leveraging a remote sensing index as a reconstruction target, are all glossed over. These kinds of adaptations for RS data inform model performance and effective downstream use, and more importantly, future research directions. We fill that gap here by discussing these unique adaptations.
@@ 3. Data for Training Remote Sensing Foundation Models
The CV field has primarily been concerned with natural images, which only sense the red, green, and blue wavelengths of the visible spectrum, and are typically captured at ground level. Some models, such as CLIP (Radford et al., 2021), FLIP (Jia et al., 2021), VisualBERT (Li et al., 2019), ViLBERT (Lu et al., 2019), FLAVA (Singh et al., 2022), and DeepSeek-VL2 (Wu et al., 2024) are VLMs that also utilize text as an additional mode of data. However, most of the RS foundation model frameworks borrowed from CV concern themselves with only the image modality. Within the RS domain, the variety of unique sensors available ranges across a number of instruments aboard different satellites. The data modes detailed below are available at mass scale and readily accessible for training foundation models. The training of RS foundation models is our focus in this manuscript, rather than providing a full catalogue of remote sensing or geospatial data. We will keep discussions focused on sensors with high resolution (e.g., less than 40m), which are also the focus of current efforts in benchmarking datasets (Lacoste et al., 2023) and training foundation models (Zhou et al., 2024).
@@ 3.1. Optical Imagery
Optical images capture electromagnetic wavelengths in the visible and near visible bands of light. These are recorded via passive sensors, which simply measure the electromagnetic radiation originating from the sun that is reflected off of objects. This means that passive sensors are more energy efficient than active sensors, but they are affected by light and weather conditions. These sensors can only capture imagery during the daytime, and can be obscured by atmospheric phenomena like cloud cover or haze. This means their use is limited in hurricanes, storms or active floods under overcast conditions.
Optical images are often acquired from different satellites. Several datasets from different satellites are openly available for free, mass download, such as MODIS (Justice et al., 2002), the Landsat (Wulder et al., 2022) family, and Sentinel-2 (Spoto et al., 2012), with the latter two offering high resolution. Each satellite contains a sensor with different spatial and spectral resolutions, but each is broadly designed to capture the visible, near-infrared, and short-wave infrared portions of the electromagnetic spectrum. In the case of the Sentinel-2 satellite, there are 13 distinct bands that are captured at resolutions ranging from 10m/pixel to 60m/pixel. These bands can be broken down into five groups based on usage as follows (Wang et al., 2016):
@@
B5-8 refer to VNIR (Visible and Near Infrared) light, and are typically used to measure vegetation health. B8a is a Narrow Near Infrared (NIR) band for moisture-based vegetation health monitoring.
B9, still in NIR spectrum, focuses on water vapor absorption and is used to measure atmospheric moisture presence. It is paired with B8A to derive total precipitable water.
B10-B11 are SWIR (Short Wave Infrared) bands. B10 is used for cirrus cloud detection. B11 is used for soil and vegetation moisture detection and snow/ice vs. cloud differentiation. B12 is often used with B2 to measure geological features and map burned areas.
This range of spectrum and information captured by Sentinel-2 is an example of the wealth of information captured by remote sensing optical products that go beyond RGB. However, CV models are designed to operate with three input channels of RGB of ‘natural’ images (obtained by regular cameras and widely available on the internet). In order to leverage the highly sophisticated pre-trained weights that are available for CV backbones, such as the ubiquitous ResNet50 (He et al., 2015) or ViT16 (Dosovitskiy et al., 2021), some RS models only utilize the visible bands of light from optical multispectral (MS) satellite images (Mañas et al., 2021; Chen et al., 2023). Some RS models also rely on RGB images taken at ground level (Cepeda et al., 2023), or from some sort of aerial platform, such as a drone or helicopter (Liu et al., 2024). While RS models trained purely on RGB can achieve SOTA performance in certain settings (Mañas et al., 2021), deep learning models in RS that only train on RGB data have an information disadvantage compared to models that train on all bands. RGB-only models can struggle to perform on a number of downstream tasks in RS as a result (Lacoste et al., 2023). The spectral range of interest and the specific choice of bands on a satellite are the result of decades of experience, in-lab experiments and consensus-building efforts to ensure satellites are equipped with bands that can assist in target-specific tasks. Figure 1 highlights the information lost when solely examining RGB imagery. For instance, infrared bands are crucial for monitoring vegetation (Neinavaz et al., 2021), as seen in figure 1(f), while NIR bands in confluence with blue and green bands can help with water quality analysis (Cillero Castro et al., 2020). SWIR can be used for measuring soil and vegetation moisture, and is depicted in figure 1(e).
While an increasing number of foundation models leverage most bands present in optical imagery such as Sentinel-2, they typically do so by a simple manipulation of input channel dimensions to the image encoder of the model (Wang et al., 2022b). The optimal architecture required for remote sensing architectures is not well-defined yet (Rolf et al., 2024).
@@ 3.2. Synthetic Aperture Radar
Synthetic Aperture Radar (SAR) is an active remote sensor that emits energy in microwave frequencies towards the earth and then captures the amount of energy reflected back. Capturing the image from a moving sensor (in this case, a satellite) allows for capturing the image through a large ’synthetic’ aperture, compared to the much smaller physical sensor aperture. This results in a much higher resolution than other sensors that monitor similar frequencies (Brown and Porcello, 1969). Each SAR sensor operates in a specific wavelength, referred to as bands, which serve different uses (Tsokas et al., 2022). Larger wavelengths provide more penetration through obstruction, such as vegetation, but result in lower resolution. One of the most commonly used bands is C-band (7.5-3.8 cm wavelength) (Tsokas et al., 2022), which is currently captured by Sentinel-1 (Torres et al., 2012), RADARSAT-1 (Mahmood, 2014), RADARSAT-2 (Morena et al., 2004), RADARSAT Constellation Mission (RCM) (Thompson*, 2015) satellites. C-band imagery is made available publicly, and is useful for global mapping and change detection. The larger wavelength provides moderate penetration still at relatively high spatial resolution (e.g., 40m). However, there are larger wavelengths available, such as L-band, which is captured by AirSAR (Hoekman and Bouman, 1993), PALSAR (Rosenqvist et al., 2007), and the soon-to-be-launched NISAR (Chapman and Rosen, 2024). L-band is often used for biomass and vegetation mapping (Haldar et al., 2012). There is also X-band, whose wavelength is smaller than C-band and often used for high resolution monitoring of fine surface features, such as vegetation structure, and urban infrastructure. (Solari et al., 2017). X-band is obtained by platforms such as Capella (Castelletti et al., 2021), ICEYE (Ignatenko et al., 2020), TerraSAR-X (Werninghaus and Buckreuss, 2010), and TanDEM-X (Krieger et al., 2007). Different wavelengths of SAR all capture different information, and different wavelengths often complement each other if available for the same region and time.
@@
SAR has another characteristic distinct from optical imagery: SAR takes advantage of materials’ innate dielectric properties. Every material has a dielectric constant that describes how strongly it interacts with electromagnetic waves. Materials with a high dielectric constant, such as water, reflect microwave energy back more strongly, allowing it to be picked up more easily by SAR (Barowski et al., 2019). This can prove useful for applications such as mapping waterways in a region that might be otherwise undetected by optical sensors. As seen in figure 1, SAR imagery clearly highlights water features that optical sensors struggle to pick up. SAR is an all-sky sensor, meaning that it operates regardless of light conditions. Its wavelength range is also large enough that it penetrates cloud cover. This combination of traits helps provide continuous spatial coverage in geographic areas that might otherwise be occluded to optical satellite sensors due to cloud cover (Brown and Porcello, 1969), or polar regions that lack sunlight for half of a year.
@@
SAR is predated by passive microwave sensing, which monitors similar microwave frequencies in order to pierce cloud cover but does not emit a signal of its own in order to do so. As a result, passive microwave sensing offers much lower resolution compared to SAR. Sensors such as AMSR (Tachi et al., 1989), AMSR-2 (Alsweiss et al., 2015), SMAP (Entekhabi et al., 2010), and GPM (Munchak et al., 2020) all passively monitor microwave signals emitted from the Earth. By comparison, SAR actively beams energy down to Earth’s surface and records the response, which allows for multiple polarizations at transmission and reception. Polarization describes the orientation of the plane that the electromagnetic wave oscillates on. The shorthand for polarization is the orientation of the emission, which is either vertical (V) or horizontal (H), followed by the orientation of the reception, which will also be either vertical or horizontal. Each polarization results in the waves scattering differently upon reflection, enabling certain polarizations to be particularly effective at picking up certain types of surface features. For example, the VV polarization is good at picking up rough surface scattering, typically caused by roughened water or vertical surfaces. HV or VH polarizations are good at picking up volume scattering, where the signal bounces around a number of times in a given volume before being broadcast back out to space. This is typically caused by forest canopies and vegetation. We see in figure 1(a) and figure 1(b) that the VH polarization provides more fine-grained details in a variety of forested areas than the VV polarization, where those areas are whited out. HH polarization is good at picking up horizontal surfaces, such as smooth ice or calm water.
@@
Compared to optical imagery, SAR captures distinct and complementary information, despite it also being stored as a raster (gridded) data type. The different frequencies of SAR, the different polarizations, and the volume scattering aspect make it stand apart from optical. For instance, surface reflectance of sea ice will appear as a bright surface in optical imagery, regardless of its age (which is a proxy for its thickness and the level of hazard it poses to marine navigation). However, as sea ice ages (from thinner first-year ice to thicker multi-year ice), its salinity changes. Because of SAR’s volume scattering (caused by salt content and air bubbles in ice), the appearance of thick ice changes in SAR imagery compared to younger, thinner ice (Pires de Lima et al., 2023). This is the primary reason sea ice analysts rely on SAR imagery in making sea ice charts and use optical only as a secondary reference. As we demonstrate later in this manuscript, this complementary aspect of SAR imagery (despite it being stored as an image) motivates incorporating this multi-sensor information into RS foundation models.